In the following memory dump file there is a 6 wor
Answer #1

Given image contains the memory dump from the address of CS:270 and here each byte holds two digits (4 bits +4 bits). Now considering the value at address CS:0286 = 14 and from here we need to consider 6 Words and as we already know that 1 Word = 2 Bytes and so 6 Words includes all the location values from

CS:0286, CS:0287, CS:0288, CS:0289, CS:028A, CS:028B, CS:028C, CS:028D, CS:028E, CS:028F, CS:0290, CS:0291

=> 14 00 1D 00 F2 FF 23 00 FC FF 1B 00

These values which are considered as 6 words (6 *2 = 12 Bytes) from the address CS:0280 and now these values are to be converted to decimal from the hexadecimal notation..

Converting hexadecimal number to decimal:

consider XY is the number represented in hexadecimal then to convert it into decimal first write the 4-bit representation for each digit and then convert the obtained binary value into decimal by multiplying with powers of 2

CS:0286 value= 14 in hexadecimal notation

0001 0100 in binary notation

2^7*0+2^6*0+2^5 *0+2^4 *1 +2^3 *0+ 2^2 *1+2^1 *0+2^0 *0

=16+4=20 in decimal notation

CS:0287 value=00 in hexadecimal notation

0000 0000 in binary

00 in decimal notation

CS:0288 value=1D in hexadecimal notation

0001 1101 in binary notation

  2^7*0+2^6*0+2^5 *0+2^4 *1 +2^3 *1+ 2^2 *1+2^1 *0+2^0 *1

=29

CS:0289 value =00 in hexadecimal notation

0000 0000 in binary notation

00 in decimal notation

Cs:028A value=F2 in hexadecimal notation

11110010 in binary notation

  2^7*1+2^6*1+2^5 *1+2^4 *1 +2^3 *0+ 2^2 *0+2^1 *1+2^0 *0

242 in decimal notation

CS:028B value=FF in hexadecimal notation

1111 1111 in binary notation

  2^7 *1+2^60 *1+2^5 *1+2^4 *1 +2^3 *1+ 2^2 *1+2^1 *0+2^0 *1

255 in decimal notation

CS:028C value=23 in hexadecimal notation

0010 0011 in binary notation

  2^7*0+2^6*0+2^5 *1+2^4 *0 +2^3 *0+ 2^2 *0+2^1 *1+2^0 *1

35 in decimal notation

CS:028D value=00 in hexadecimal

0000 0000 in binary

00 in decimal

CS:028E value=FC in hexadecimal

1111 1100 in binary

  2^7*1+2^6*1+2^5 *1+2^4 *1 +2^3 *1+ 2^2 *1+2^1 *0+2^0 *0

252 in decimal

CS:028F value=FF in hexadecimal

1111 1111 in binary

2^7*1+2^6*1+2^5 *1+2^4 *1 +2^3 *1+ 2^2 *1+2^1 *1+2^0 *1

255 in decimal

CS:0290 value= 1B in hexadecimal

0001 1011 in binary

2^7*0+2^6*0+2^5 *0+2^4 *1 +2^3 *1+ 2^2 *0+2^1 *1+2^0 *1

27 in decimal

Cs:0291 value=00 in hexadecimal

0000 0000 in binary

00 in decimal

So all the 6 words from address location CS:0286 are converted into decimal values